# TelePreview: A User-Friendly Teleoperation System with Virtual Arm Assistance for Enhanced Effectiveness

### v4

@@ Abstract
Teleoperation provides an effective way to collect robot data, which is crucial for learning from demonstrations. In this field, teleoperation faces several key challenges: user-friendliness for new users, safety assurance, and transferability across different platforms. While collecting real robot dexterous manipulation data by teleoperation to train robots has shown impressive results on diverse tasks, due to the morphological differences between human and robot hands, it is not only hard for new users to understand the action mapping but also raises potential safety concerns during operation. To address these limitations, we introduce TelePreview. This teleoperation system offers real-time visual feedback on robot actions based on human user inputs, with a total hardware cost of less than \$1,000. TelePreview allows the user to see a virtual robot that represents the outcome of the user’s next movement. By enabling flexible switching between command visualization and actual execution, this system helps new users learn how to demonstrate quickly and safely. We demonstrate that it outperforms other teleoperation systems across five tasks, emphasize its ease of use, and highlight its straightforward deployment across diverse robotic platforms. We release our code and a deployment document on our website https://nus-lins-lab.github.io/telepreview/.
@@ I Introduction
Figure 2: Overview of the TelePreview System Architecture: Our system consists of three main components: (1) Various input devices, including RGB-D cameras, IMU mocap suits, VR headsets, mocap gloves, AR hand tracking, and EMG armbands for capturing human motion; (2) A central processing pipeline that uses SMPL-X as a standard to process wrist and hand pose data, feeding into our Preview System with joint-to-joint mapping and a non-collision retargeting network; (3) Support for different robot platforms as output devices for executing the mapped movements.
Researchers have long recognized teleoperation as an essential component for gathering on-robot data used in learning from demonstrations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. An intuitive and user-friendly teleoperation system is crucial for collecting high-quality, diverse, and scalable data. However, effectively controlling robots with dexterous hands remains a significant challenge. The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex. These inherent structural and functional disparities, combined with the precision required for fine-grained tasks, often result in inefficient and potentially compromised data collection processes.
@@ II Related Work
Figure 3: Pipeline of Our Modules: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides visual preview before physical execution. We achieve precise alignment between virtual and physical robots through AprilTag calibration.
II-A Robot Teleoperation Frameworks for Manipulation
The growing interest in training robots through imitation learning has driven the need for extensive, high-quality real robot datasets. Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks[13, 14, 15]. Researchers have explored various teleoperation systems, including VR controllers [1, 2, 3], motion capture [16, 17, 18], wearable gloves [19, 11, 20, 21], exoskeletons [22, 23], each offering distinct benefits in terms of accessibility, precision, and generalizability.
Vision-based teleoperation systems [6] typically employ RGB or RGB-D cameras to detect hand poses. The hand pose retargeting module maps the human hand pose data obtained from perception algorithms into joint positions of the dexterous hand. We formulate this process as an optimization problem that minimizes the difference between the key point vectors of the human hand and the dexterous hand. To generate the end-effector motions, 3D positions of key points in the camera frame are calculated by using an RGB-D camera. The wrist pose is then computed by keypoint positions in the local wrist frame and global camera frame using the Perspective-n-Point (PnP) algorithm.
Although vision-based teleoperation systems offer a relatively low-cost solution [6, 2, 3, 1], they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspace, and high computational demands. This limitation stems from the requirement of vision-based approaches for operators to keep their hands within the camera’s field of view, thereby restricting the action space of the human hand. Although increasing the scale factor can amplify end-effector movement and expand the action space of the robot, it aggravates the difficulty for operators in performing fine-grained tasks. These approaches also struggle to handle scenarios where fingers overlap. The extra keypoint retargeting module can be computationally expensive in real-time. Although specialized hardware [4] can improve accuracy, it substantially increases system costs. Our proposed system directly uses the joint pose obtained from the motion capture suit and data gloves, preventing the deviation of vision-based systems and reducing the computational cost. Recent unified teleoperation frameworks attempt to address these limitations but often introduce new trade-offs. Some approaches [3, 6] prioritize implementation simplicity by tracking only fingertip positions rather than enabling detailed joint-level control. Furthermore, your customization options typically require time-consuming manual parameter tuning that requires significant expertise.
@@ II-B Visual Feedback in Teleoperation Systems
To address these limitations, researchers have explored enhanced visualization techniques [26, 27, 28, 29, 30] that overlay virtual information on the workspace view. Even though these approaches improve user perception by displaying planned trajectories and environmental constraints, they primarily design and visualize AR objects for a specific task, such as a pink box with dimensions of 10 cm × 10 cm × 10 cm. Any changes to the object’s attributes (e.g., dimensions, color, or type) invalidate the preconfigured AR object, so the system must rebuild it to match the new characteristics.
Instead of visualizing the manipulated objects, our system provides an AR representation of the robot that directly maps operator commands. This design ensures safety for new users who may be unfamiliar with how the robot responds to their inputs. Our configuration allows users to use their own URDF file, eliminating the need to reconstruct manipulated objects for every new task.
@@ III System Overview
    Once the I/O state transitions from active to inactive, the final preview pose is captured and converted into an optimal trajectory for actual execution (Section V-B3).
@@ IV-B Hand Pose Estimation
Mocap gloves [33] offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors. Our joint-to-joint mapping approach provides a generalizable solution for retargeting high-dimensional human hand motion data to lower-dimensional robotic systems. This approach is particularly effective when the input space (e.g., the human hand with 27 DoF) has higher dimensionality than the target space (e.g., the robotic hand with 16 DoF), as it allows selective mapping of the most intuitive and task-relevant degrees of freedom. Through manual selection of corresponding joints based on human intuition and task requirements, followed by range alignment between input and output spaces, we maintain natural and predictable motion correspondence.
@@ IV-C Non-Collision Retargeting Network
Direct mapping from human hand postures to robotic hands using existing retargeting approaches like [6] can result in self-collisions, as demonstrated in Fig. 4(a). While traditional retargeting methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations.
@@
As demonstrated in Fig. 4(b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz).
@@ V-A Why Preview System Assistance?
Traditional teleoperation approaches face significant challenges in data quality when collecting demonstrations for imitation learning. During complex manipulation tasks, users frequently make exploratory movements that contaminate the demonstration data with sub-optimal trajectories. These repeated probing actions to locate the socket edge and refine the approach angle before successful insertion introduced significant noise into the collected datasets. Such noise is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements.
@@ V-B Technical Implementation of Preview System
We design our preview feature to integrate seamlessly with the teleoperation workflow. It requires three critical elements: (1) precise spatial calibration to maintain alignment between the virtual preview and the physical robot, (2) realistic rendering to blend the preview with camera feeds, and (3) state management to handle transitions between preview and execution phases. The following subsections detail each of these technical components, highlighting how they collectively enable users to view and adjust planned motions prior to committing them to the physical robot.
@@ V-B1 Spatial Alignment Using AprilTag
To maintain a precise spatial alignment between the preview and the physical robot in the camera viewpoints, we implement a robust calibration system using AprilTags [31]. Our approach involves two sequential steps: (1) a hand-eye calibration that establishes the transformation between the robot’s base frame and the fixed camera frame, and (2) AprilTag pose detection that enables real-time tracking of the robot’s position and orientation relative to each camera view. This chain of transformations ensures consistent preview visualization across different viewpoints.
@@
As shown in Fig. 5, our calibration procedure involves multiple sequential transformations. We begin with hand-eye calibration of the fixed top-view camera (1), followed by detecting the AprilTag’s pose from this camera’s perspective (2). Using these relationships, we can compute the transformation between the AprilTag and the robot base frame (2). For the additional floating third-view camera, we first detect the AprilTag pose from its viewpoint (3), which then allows us to establish its position relative to the robot base frame through the standard AprilTag reference (3). This chain of transformations ensures consistent preview visualization regardless of camera movement or viewpoint changes, making our system robust for dynamic viewing scenarios.
@@ V-B2 Visual Preview Visualization
Building upon this spatial alignment system, we implement the visual component of our preview system. As shown in Fig. 6(a), we use the pyrender [35] library to render the virtual robot through the previous float camera. The rendered preview image is then composited with the actual camera feed using alpha blending, creating a seamless overlay where the virtual robot appears naturally integrated with the physical environment. This visualization approach requires careful calibration of the virtual camera parameters to match the physical cameras, ensuring that the preview accurately reflects the robot’s intended configuration in the workspace. The resulting overlay provides users with an intuitive preview of planned motions. Our preview system also supports multiple viewpoints to enhance spatial awareness during complex manipulation tasks. As shown in Fig. 6(b) and Fig. 6(c), we deploy cameras to capture both third-person and top-down views of the workspace, allowing users to verify planned movements from complementary perspectives.
@@
TABLE I: Real Robot Teleoperation Results. We use the success rates the baseline methods report in their papers, marking tasks they did not attempt or do not support with “-”. Success rates for Open Teach [3] reflect expert performance.
(a) Preview Rendering
(b) Third-person View
(c) Top-down View
Figure 6: Rendering and Multi-view Visualization System.
@@
Pick & Place 0.6 1.0 +0.4 23.56 13.55 -10.01
Hang 0.6 1.0 +0.4 29.30 30.83 +1.53
Pour 0.9 0.8 -0.1 43.20 36.13 -7.07
Box Rotation 0.6 0.8 +0.2 30.47 19.12 -11.35
Cup Stacking 0.5 1.0 +0.5 31.54 18.91 -12.63
@@
V-B3 IO Control for Preview Visualization
Our system uses a foot pedal as an intuitive control interface for state transitions. As shown in Fig. 7, when the IO is activated, the physical robot stops while the preview appears and responds to user commands for motion preview (Preview Mode). Upon IO deactivation, the preview disappears, and the system extracts its final configuration as the target pose to align, bypassing any exploratory movements made during the preview (Align Mode). This target is then processed by the mplib motion planning library [37] to generate an optimal trajectory for execution. After reaching the target pose, the robot automatically resumes real-time teleoperation until the next IO is activated. This workflow ensures users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
Figure 7: State Transition of Preview Control
@@ VI-A Experiment Setup
TelePreview integrates several key components. We capture the user’s movements via an inertial motion capture suit [38] and a pair of strain-gauge-based mocap gloves [33], providing high-fidelity joint angle data for both body posture and delicate hand movements. To visualize robot action before execution, we employ a RealSense RGB-D camera and render the preview either on a standard 2D display or within a VR headset (e.g., Meta Quest 3[39]). For the physical hardware, we use a UFactory xArm-6 robotic arm[40] outfitted with a Leap hand [34], utilizing AprilTag [31] to align the initial poses of the preview and the real robot.
@@ VI-B Task Descriptions
We test TelePreview on five real-world tasks, each highlighting different manipulation challenges:
@@ VI-C Evaluation Metrics
  • Success Rate: We define the success rate as the proportion of times the operator completes the task within the given constraints. For example, if the user attempts a cup-hanging task $X$ times and successfully hangs the cup on the rack $Y$ times, the success rate is $\frac{Y}{X}$.
@@ VI-D Performance Comparison with Baselines (Q1)
To address Q1—evaluating how our approach compares to existing teleoperation methods—we measure the actual robot performance of TelePreview against three baselines: Telekinesis [36], AnyTeleop [6], and Open Teach [3]. For fairness and consistency, an expert user operates each method and tests it on the five tasks described in Section VI-B. The primary metrics are success rate and average execution time.
As summarized in Table I, TelePreview achieves higher success rates than all baselines. While Telekinesis, AnyTeleop, and Open Teach handle simpler, short-horizon tasks, they often struggle with complex manipulations requiring subtle wrist rotations and fine-grained positioning (e.g., the Hang task). TelePreview, by contrast, captures these nuances more effectively, reflecting stronger overall performance and versatility. These findings demonstrate TelePreview’s ability to address a broad spectrum of real-world challenges, from basic pick-and-place operations to more intricate tasks involving delicate movements.
(a) LEAP Hand
(b) Gripper
Figure 8: Deployment on Different Robots.
@@ VI-E Effect of Preview for Users (Q2)
To investigate Q2—how the preview feature benefits new (non-expert) users—we evaluate both success rate and average execution time in two conditions: (1) with the preview enabled and (2) with the preview disabled. The same five tasks from Section VI-B are attempted by new users, with each task repeated 10 times.
As shown in Table II, enabling the preview boosts user performance. Success rates increase across all tasks, with especially notable gains in challenging manipulations such as cup stacking (+0.5) and pick & place (+0.4). Moreover, average execution times drop by up to 12.63 seconds (cup stacking), demonstrating not only improved accuracy but also greater efficiency. By visualizing a virtual “preview” of the robot’s intended motion and confirming the final pose prior to executing physical movements, users eliminate the need for exploratory motions in the actual workspace. This design ensures that recorded trajectories contain only intentional, task-oriented actions without extraneous trial-and-error, providing cleaner, high-quality demonstrations for imitation learning. As a result, TelePreview delivers both a more intuitive teleoperation experience and more consistent data for downstream training.

### v6

@@ Abstract
Teleoperation provides an effective way to collect robot data, which is crucial for learning from demonstrations. In this field, teleoperation faces several key challenges: user-friendliness for new users, safety assurance, and transferability across different platforms. While collecting real robot dexterous manipulation data by teleoperation to train robots has shown impressive results on diverse tasks, due to the morphological differences between human and robot hands, it is not only hard for new users to understand the action mapping but also raises potential safety concerns during operation. To address these limitations, we introduce TelePreview. This teleoperation system offers real-time visual feedback on robot actions based on human user inputs, with a total hardware cost of less than \$1,000. TelePreview allows the user to see a virtual robot that represents the outcome of the user’s next movement. By enabling flexible switching between command visualization and actual execution, this system helps new users learn how to demonstrate quickly and safely. We demonstrate that it outperforms other teleoperation systems across five tasks, emphasize its ease of use, and highlight its straightforward deployment across diverse robotic platforms. We release our code and a deployment document on our website https://nus-lins-lab.github.io/telepreview-web/
@@ I Introduction
Figure 2: Overview of the System Architecture (1) Various input devices for capturing human motion. Only one device from each of the two input groups (wrist pose and hand gesture) is required during operation. The VR headset is shown at the intersection of the two groups, as it can capture both; (2) A processing pipeline based on SMPL-X that performs joint mapping and collision-free retargeting; (3) Output to various robot platforms for executing the mapped motions.
Researchers have long recognized teleoperation as an essential component for gathering on-robot data used in learning from demonstrations [1, 2, 3, 4, 5, 6]. An intuitive and user-friendly teleoperation system is crucial for collecting high-quality, diverse, and scalable data. However, effectively controlling robots with dexterous hands remains a significant challenge. The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex [7, 8, 9]. These inherent structural and functional disparities, combined with the precision required for fine-grained tasks, often result in inefficient and potentially compromised data collection processes.
@@ II Related Work
Figure 3: Pipeline of Our Modules: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides a visual preview before physical execution. We achieve precise alignment between virtual and physical robots through AprilTag calibration.
II-A Teleoperation Frameworks for Dexterous Manipulation
Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks[10, 11, 12]. But Many existing teleoperation frameworks focus on low-degree-of-freedom (low-DoF) robotic grippers, using input modalities such as keyboards[13], joysticks[14, 15], or gamepads[16] to issue discrete commands. While these are effective for simple tasks, they are not sufficient for the fine-grained control of dexterous robotic hands with High DoFs.
Researchers have explored various teleoperation systems for dexterous manipulation, including VR headsets [7, 17, 1], motion capture [18, 19, 20], wearable gloves [21, 6, 22, 23, 24], exoskeletons [25, 26].
Using VR headsets, vision-based teleoperation systems [3, 17, 1, 7] typically employ RGB cameras to detect hand poses. These systems often rely on estimated key points from visual perception to map user motions onto the robot, typically using optimization-based retargeting methods.
Although vision-based teleoperation systems offer a relatively low-cost solution, they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspaces, and high computational demands. These systems also struggle with occlusions, finger overlap, and other perception issues. The keypoint-based retargeting step can be computationally expensive in real time. Although specialized hardware [2] can improve accuracy, it substantially increases system costs.
In contrast, exoskeleton systems[25, 26] provide highly accurate motion capture with low latency but tend to be tightly coupled to specific hardware platforms, which limits their generalizability across different robot embodiments.
In this work, we adopt a glove-based teleoperation framework that captures human hand motion at the joint level and maps it to a unified body model (e.g., SMPL-X) before retargeting it to the robot. This abstraction layer enables real-time, joint-level control of dexterous robotic hands while improving generalizability across different robot embodiments. Although glove-based teleoperation has previously been considered expensive, recent commercialization and the availability of low-cost, off-the-shelf data gloves have made this approach increasingly accessible.
@@ II-B Visual Feedback in Teleoperation Systems
To address these limitations, researchers have explored enhanced visualization techniques [29, 30, 31, 32] in AR-based teleoperation systems[33, 34], which aim to improve user understanding of spatial constraints and planned motions.
Prior AR-based teleoperation systems typically display a virtual robot arm in the environment but do not spatially align it with the physical robot. As a result, they cannot capture interaction data during contact with real objects or the robot’s behavior following physical execution. In contrast, our system overlays the AR robot directly on top of the physical robot, enabling users to first control a virtual preview and then switch to executing the same motion with the real hardware. This design allows us to capture the full sequence of a teleoperation task—from intent to contact to post-contact behavior—without requiring real-to-sim translation or object re-modeling. By eliminating the need to reconstruct physical objects in simulation, we reduce system setup complexity while enabling high-fidelity, real-world demonstration collection.
@@ III System Overview
    Once the I/O state transitions, the final preview pose is captured and converted into an optimal trajectory for actual execution (Section V-B3).
@@ IV-B Hand Pose Estimation
Mocap gloves offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors. Our joint-to-joint mapping approach provides a generalizable solution for retargeting high-dimensional human hand motion data to lower-dimensional robotic systems. This approach is particularly effective when the input space (e.g., the human hand with 27 DoF) has higher dimensionality than the target space (e.g., the robotic hand with 16 DoF), as it allows selective mapping of the most intuitive and task-relevant degrees of freedom. Through a manual selection of corresponding joints based on human intuition and task requirements, followed by range alignment between input and output spaces, we maintain natural and predictable motion correspondence.
@@ IV-C Non-Collision Retargeting Network
Direct mapping from human hand postures to robotic hands using existing retargeting approaches like [3] can result in self-collisions, as demonstrated in Fig. 4(a). While traditional retargeting methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations, highlighting the importance of collision-aware methods such as [32].
@@
As demonstrated in Fig. 4(b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz). The implementation detail can be found in the Appendix.
@@ V-A Why Preview System Assistance?
Traditional teleoperation approaches face significant challenges in data quality when collecting demonstrations for imitation learning. During complex manipulation tasks, users frequently make exploratory movements that contaminate the demonstration data with sub-optimal trajectories. Such noise is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements. Recent research has highlighted how errors or exploratory movements in teleoperation can decrease the generalization ability of the learned policies and negatively impact performance on unseen tasks.[38, 39, 40, 41, 42]
@@ V-B Technical Implementation of Preview System
We designed our preview feature to integrate seamlessly with the teleoperation workflow. It requires three critical elements: (1) precise spatial calibration to maintain alignment between the virtual preview and the physical robot, (2) realistic rendering to blend the preview with camera feeds, and (3) state management to handle transitions between preview and execution phases. The following subsections detail each of these technical components, highlighting how they collectively enable users to view and adjust planned motions prior to committing them to the physical robot.
@@ V-B1 Spatial Alignment Using AprilTag
To maintain a precise spatial alignment between the preview and the physical robot in the camera viewpoints, we implement a robust calibration system using AprilTags [35]. Our approach involves two sequential steps: (1) a hand-eye calibration that establishes the transformation between the robot’s base frame and the fixed camera frame, and (2) AprilTag pose detection that enables real-time tracking of the robot’s position and orientation relative to each camera view.
@@
As shown in Fig. 5, our calibration procedure involves multiple sequential transformations. We begin with hand-eye calibration of the fixed top-view camera, followed by AprilTag pose detection from this camera’s perspective. These steps allow us to compute the transformation from the AprilTag to the robot base frame. For the floating third-view camera, we detect the AprilTag pose and use it to determine its position relative to the robot base frame. This calibration ensures consistent visualization across dynamic or multi-view setups.
@@ V-B2 Visual Preview Visualization
Building upon this spatial alignment system, we implement the visual component of our preview system. As shown in Fig. 6(a), we use the pyrender library to render the virtual robot through the float camera. The preview image is then composited with the real camera feed using alpha blending, creating an overlay where the virtual robot appears aligned with the physical environment. This requires the calibration of virtual camera parameters to match the physical cameras, ensuring the preview accurately reflects the robot’s intended configuration. The preview offers real-time visual guidance before execution. Our system supports multiple viewpoints to aid spatial reasoning and trajectory checking. As shown in Fig. 6(b) and Fig. 6(c), we deploy third-person and top-down cameras, rendered and shown simultaneously on an external monitor. The system does not require a VR headset. For instance, during manipulation tasks, users can simultaneously monitor the robot’s vertical position from the third-person view while checking its planar alignment from the top-down view.
V-B3 IO Control for Preview Visualization
Our system uses a foot pedal as an intuitive control interface for state transitions. As shown in Fig. 7, when the IO is activated, the physical robot stops while the preview appears and responds to user commands for motion preview (Preview Mode). Upon IO deactivation, the preview disappears, and the system extracts its final configuration as the target pose to align, bypassing any exploratory movements made during the preview (Align Mode). The mplib motion planning library then processes this target to generate an optimal trajectory for execution. During this motion, the user does not need to remain static—the robot executes the planned motion autonomously. Once the target pose is reached, the system enters a continuous teleoperation mode in which the robot follows the user’s live movements in real time until the next preview is initiated. This workflow ensures that users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
(a) Preview Rendering
(b) Third-person View
(c) Top-down View
Figure 6: Rendering and Multi-view Visualization System.
Figure 7: State Transition of Preview Control
@@ V-B3 IO Control for Preview Visualization
TABLE I: Real Robot Teleoperation Results We mark tasks they did not attempt or do not support with “-”. Success rates for Open Teach [1] reflect expert performance.
@@
Pick & Place 0.6 1.0 +0.4 23.56±$\pm 4.65$ 13.55±$\pm 3.17$10.01
Hang 0.6 1.0 +0.4 29.30±$\pm 8.18$ 30.83±$\pm 3.65$ +1.53
Pour 0.9 0.8 -0.1 43.20±$\pm 6.73$ 36.13±$\pm 5.86$7.07
Box Rotation 0.6 0.8 +0.2 30.47±$\pm 14.08$ 19.12±$\pm 3.32$11.35
Cup Stacking 0.5 1.0 +0.5 31.54±$\pm 6.04$ 18.91±$\pm 1.29$12.63
@@ VI-A Experiment Setup
TelePreview integrates several key components. We capture the user’s movements via an inertial motion capture suit and a pair of strain-gauge-based mocap gloves, providing high-fidelity joint angle data for both body posture and delicate hand movements. To visualize robot action before execution, we employ a RealSense RGB-D camera and render the preview either on a standard 2D display or within a VR headset (e.g., Meta Quest 3). For the physical hardware, we use a UFactory xArm-6 robotic arm outfitted with a Leap hand [37], utilizing AprilTag [35] to align the initial poses of the preview and the real robot.
@@ VI-B Task Descriptions
We test TelePreview on five real-world tasks. These tasks were carefully selected to cover a diverse range of manipulation skills, including grasping, rotation, alignment, and insertion, and are representative of common benchmarks used in teleoperation research:
@@ VI-C Evaluation Metrics
  • Success Rate: We define the success rate as the proportion of times the operator completes the task within the given constraints.
@@ VI-D Performance Comparison with Baselines (Q1)
Due to frequent self-collision failures encountered during local deployment of vision-based methods, we report results from their original publications and discuss reproduction experiences in the Appendix. As summarized in Table I, TelePreview achieves higher success rates than all baselines. While Telekinesis, AnyTeleop, and Open Teach handle simpler, short-horizon tasks, they often struggle with complex manipulations requiring subtle wrist rotations and fine-grained positioning (e.g., the Hang task). TelePreview, by contrast, captures these nuances more effectively, reflecting stronger overall performance and versatility.
@@ VI-E Effect of Preview for Users (Q2)
To investigate Q2—how the preview feature benefits new (non-expert) users—we evaluate both success rate and average execution time in two conditions: (1) with the preview enabled and (2) with the preview disabled. Our study involved 10 participants in total, including 5 new users with no prior experience in robotic teleoperation and 5 experienced users (3 of whom are authors). This section focuses on the new user group.
Participants were given consistent task instructions and evaluated under a standardized protocol, including predefined success criteria. Task order was counterbalanced across participants to mitigate ordering effects.
The same five tasks from Section VI-B are attempted by new users, with each task repeated 10 times. We report both mean and standard deviation across trials for each task. In addition to objective performance, we collected subjective workload ratings using a 5-point Likert scale (using the same evaluation metrics in [43]) and recorded the practice time required for participants to reach task readiness under different input modalities. Detailed results and analysis are provided in the Appendix.
As shown in Table II, enabling the preview boosts user performance. Success rates increase across all tasks, with especially notable gains in challenging manipulations such as cup stacking (+0.5) and pick & place (+0.4). Moreover, average execution times drop by up to 12.63 seconds (cup stacking), demonstrating not only improved accuracy but also greater efficiency. The standard deviations across trials are also reduced, suggesting more consistent user behavior.
By visualizing a virtual “preview” of the robot’s intended motion and confirming the final pose prior to executing physical movements, users eliminate the need for exploratory motions in the actual workspace. This design ensures that recorded trajectories contain only intentional, task-oriented actions . This results in higher-quality demonstrations that are free from unnecessary trial-and-error behaviors.